OpenStudy (anonymous):

WUT Find the x- and y- intercepts for the graph of the following equation. 5x + 3y = -30

OpenStudy (anonymous):

To do this you just put 0 in for x to find the y and then after you found it go back to the problem and put 0 in for y to get the x

OpenStudy (anonymous):

x=(-6,0) y=(0,-10)

OpenStudy (anonymous):

To find the x-intercept, Let y = 0 5x + 3y = -30 5x + 3(0) = -30 5x = -30 -- ---- 5 5 x = -6 (-6, 0) To find the y-intercept, Let x = 0 5x + 3y = -30 5(0) + 3y = -30 3y = -30 --- ---- 3 3 y = -10 (0, -10)